Jennifer is 1 year older than lauren. Three times lauren's age in 2 years is the same as my mother's age 6 years ago. If my mother is 4 years younger than my father, how old will i be next year?
Answer by ankor@dixie-net.com(22740)

let L = Lauren's present age
let J = Jennifer's age
let m = mom's present age
let f = Dad's age
:
"Jennifer is 1 year older than lauren.
j = L + 1
:
"Three times lauren's age in 2 years is the same as my mother's age 6 years ago."
3(L+2) = m - 6
:
If my mother is 4 years younger than my father,
m = f - 4
:
how old will i be next year? You will be one year older